OpenStudy (anonymous):

If f(x) = |(x2 - 10)(x2 + 1)|, how many numbers in the interval 0 ≤ x ≤ 2.5 satisfy the conclusion of the mean value theorem?

OpenStudy (anonymous):

I assume they are asking you how many points the derivative is equal to the average rate of change.

OpenStudy (anonymous):

First find: \[ \frac{f(2.5)-f(0)}{2.5-0} \]

OpenStudy (anonymous):

Then find all \(x^*\) such that: \[ f'(x^*) = \frac{f(2.5)-f(0)}{2.5-0} \]

OpenStudy (anonymous):

Remember that: \[ \frac{d}{dx}|x| = \frac{|x|}{x} \]By the chain rule: \[ |f(x)|' = \frac{|f(x)|f'(x)}{f(x)} \]

OpenStudy (anonymous):

so f '(c) = 1 ?

OpenStudy (perl):

Does that say f(x) = |(x^2 - 10)(x^2 + 1)|

OpenStudy (anonymous):

yes

OpenStudy (perl):

did you follow wio's suggestion and calculate the average rate of change?

OpenStudy (anonymous):

yes and I got 1 for that

OpenStudy (freckles):

you give try to get rid the absolute value before differentiating f x^2+1 is always positive so f(x)=(x^2+1)|x^2-10| we are only conccerned what happens on x=0 to x=2.5 y=x^2-10 is a parabola that is faced up with x-intercepts -sqrt(10) and sqrt(10) with a y-intercept of -10 so between x=0 to x=2.5 the parabola is underneath the x-axis so there |x^2-10|=-(x^2-10) so I would look at differentiating f(x)=-(x^2+1) (x^2-10) but either way